OpenStudy (giovxnnx):

URGENT HELP! I know the answer for (a) so I'll write down, too. For (b) I tried "$89,000 because that is the price with the biggest unusual z-score" but I got that this part is incorrect because I needed to use the mean and standard deviation that I calculated along with z=2 to determine the threshold higher price. That got me lost. Question: Consider the following prices for a sample of 10 Congo-imported black diamonds: $34,550; $45,500; $75,000; $33,000; $89,000; $40,000; $55,000, $62,500; $88,000; $81,750. (a) Calculate mean and standard deviation for the above list of prices. Answer: Mean is 60,430 and Std Dev is 21,958.0813. (b) What is the higher price you will pay for a Congo-imported diamond based on the sample from above? Explain your rationale for that.

OpenStudy (dumbcow):

i tried to answer before. Z represents how many std dev you are from the mean \[z = \frac{x - mean}{stdev}\] \[x = mean + z*stdev\] If we use z=2 to find high price (x) \[x = mean +2*stdev\]

OpenStudy (giovxnnx):

What would be the answer and the rationale explanation for that?!